# Various Properties of Various Ultrafilters, Various Graph Width Parameters, and Various Connectivity Systems

### v6

@@ 4.2 Filter Subbase on a Connectivity System
Based on the above, the relationships shown in the following diagram become clear.
Figure 2: Relation about generating a filter
@@ 5.2 Ultrafilter on Countable Connectivity System
Game 5.19 (Ultrafilter Game on a Countable Connectivity System).
@@ 5.3 Elimination Games with an Invisible Robber on a Countable Connectivity System
Game 5.24 (Elimination Game with an Invisible Robber).
@@ 6.1.6 Examination for Directed proper-path-width
We will explore the concept of proper-path-width for graphs. Proper-path-width [655, 654, 657] measures how "tightly" a graph can be decomposed into overlapping vertex subsets, ensuring that no two subsets are identical while maintaining specific connectivity constraints. We are interested in extending this definition to directed graphs. Although still in the conceptual stage, the definition is outlined below [280].
@@ 6.1.7 Fuzzy Tree-width
Although still in the conceptual stage, the definitions are outlined below.
@@
Let $\xi=(V,\sigma,\mu,E)$ be a directed fuzzy graph. A fuzzy directed tree-decomposition of $\xi$ is a triple $(T,X,W)$, where:
@@
Let $\xi=(V,\sigma,\mu,E)$ be a directed fuzzy graph. A fuzzy directed path-decomposition of $\xi$ is a sequence $(X_{1},\dots,X_{r})$ of fuzzy subsets of $V$, called bags, that satisfies the following conditions:
@@
Let $\xi=(V,\sigma,\mu,E)$ be a directed fuzzy graph. A layout of $\xi$ is a bijection $f:V\rightarrow\{1,2,\dots,|V|\}$. The fuzzy directed bandwidth of the layout $f$ is defined as:
@@ 6.1.8 Graph parameters for Fuzzy HyperGraph
In fuzzy graph theory, the concept of a fuzzy hypergraph is well-known. It extends the notion of a hypergraph by incorporating the concepts of fuzzy sets and fuzzy graphs. Numerous studies have explored its applications in real-world scenarios[547, 61, 18, 512].
@@
We are interested in exploring whether new insights can be gained by extending the graph width parameter known as hypertree-width to fuzzy hypergraphs. Given the vast range of applications for fuzzy hypergraphs, such an extension could potentially accelerate the development of real-world applications.
@@
Let $H=(\mu,\rho)$ be a fuzzy hypergraph defined on a non-empty set $X$, where:
@@ 6.1.9 Directed Hypertree Width
We are interested in exploring whether new insights can be gained by extending the graph width parameter known as hypertree-width to directed hypergraphs. Although still in the conceptual stage, the definitions for directed hypergraph are outlined below.
Definition 6.33.
A directed hypertree decomposition of a directed hypergraph $H=(V,E)$ is a triple $D=(T,\chi,\lambda)$, where:
  • $T$ is a tree,
  • $\chi$ is a function that assigns each tree node $t\in V(T)$ a set $\chi(t)\subseteq V$, called a bag,
  • $\lambda$ is a function that assigns each tree node $t\in V(T)$ a set $\lambda(t)\subseteq E$, called an arc cover of the bag $\chi(t)$.
The decomposition must satisfy the following conditions:
  1. 1.
    Arc Cover Condition: For each hyper-arc $e\in E(H)$, there exists a tree node $t\in V(T)$ such that the vertices in $\mathcal{E}=e_{\text{Tail}}\cup e_{\text{Head}}$ are contained in the bag $\chi(t)$.
  2. 2.
    Connectedness Condition: For every vertex $v\in V(H)$, the set of tree nodes $\{t\in V(T):v\in\chi(t)\}$ induces a connected subtree of $T$.
  3. 3.
    Special Condition: For any two nodes $t,t^{\prime}\in V(T)$ such that $t^{\prime}$ is a descendant of $t$ in $T$, and for each hyper-arc $e\in\lambda(t)$, it holds that $(e\setminus\chi(t))\cap\chi(t^{\prime})=\emptyset$.
[…]
We will also explore the concept of fuzzy directed hypertree-width. Given the active research on the real-world applications of fuzzy directed hypergraphs (ex.[17, 513, 551, 553, 552]), studying fuzzy directed hypertree-width is both meaningful and timely.
@@ 6.1.10 Weighted Width
We also plan to explore width parameters such as Weighted Directed Tree-width, Weighted Directed cut-width, Weighted Directed Tree-breadth, Weighted Directed Hypertree-width and Weighted Directed Path-width in the context of weighted digraphs. Given that weighted digraphs are practical and versatile graph classes with a wide range of applications[595, 67, 548, 14, 347], defining these width parameters is both meaningful and valuable for various fields.
@@ 6.1.11 Directed Linear-width
We will consider about Directed Linear-width[282]. Although still in the conceptual stage, the definitions for weighted graph are outlined below. This is the linear version of the Branch-width of a directed graph[130] and is also an extension of the Linear-width of an undirected graph[659, 664].
@@ 6.1.12 Directed Connectivity system
The graph class known as Bidirected Graphs [110, 710] has also been a subject of study. Since it generalizes the concept of Directed Graphs, I anticipate that research in this area will continue to grow. In fact, Bidirected Tree-width has already been defined in the literature [698]. Therefore, defining parameters such as Bidirected Path-width (the path version of Bidirected Tree-width), Bidirected Cut-width, Bidirected Branch-width, Bidirected Linear-width, and Bidirected Connectivity system and exploring their relationships, would be highly meaningful. Additionally, extending these concepts to bidirected hypergraphs is another avenue worth exploring (cf.[645]).
@@ 6.2.1 Simple Game on a Connectivity System
A simple game is a mathematical model of voting systems where a coalition of agents wins if it belongs to a specified set of “winning” coalitions, adhering to certain conditions like monotonicity (cf. [274, 545, 175, 658]). These simple games model voting power and provide a generalized interpretation of the notion of what is “majority.” This game is well-known to be closely related to ultrafilters.
@@ 6.2.5 Ultrafilter Width under Small Set Connectivity Expansion Hypothesis
In computational complexity theory, a computational hardness assumption posits that a specific problem cannot be solved efficiently. One such pivotal assumption is the Small Set Expansion Hypothesis (SSEH), introduced by Raghavendra and Steurer, which plays a critical role in understanding the hardness of approximation algorithms [600]. The SSEH asserts that it is NP-hard to differentiate between small sets of vertices in a graph with low edge expansion (few edges leaving the set) and those with high edge expansion (many edges leaving the set). This hypothesis has become fundamental in analyzing the complexity of various graph algorithms. Since its inception, SSEH has led to numerous inapproximability results, including those documented in [702, 39, 714, 137, 251, 70].

### v8

@@ 4.2 Filter Subbase on a Connectivity System
4.3 Ultrafilter Number
In set theory and various mathematical fields, the concept of the Ultrafilter Number (Cardinality) is frequently discussed (ex.[564, 1011, 239]). We plan to explore whether this can be extended to Connectivity Systems and, if so, what characteristics can be derived from such an extension.
Definition 4.17 (Ultrafilter Number on a Connectivity System).
Let $(X,f)$ be a finite connectivity system, where $X$ is a finite set and $f:2^{X}\rightarrow\mathbb{N}$ is a symmetric submodular function. The Ultrafilter Number $u(X,f)$ on this connectivity system is defined as the minimum size of a set $B\subseteq X$ such that $B$ forms a ultra prefilter for a non-principal ultrafilter on $(X,f)$. Formally,
$u(X,f)=\min\{|B|:B\text{ is a ultra prefilter for a non-principal ultrafilter on }(X,f)\}.$
Question 4.18.
How does the Ultrafilter Number relate to the order on a Connectivity System?
@@ 5.1 Ultrafilter on Infinite Connectivity System
5.1.1 P-point and Q-point on an infinite connectivity system
A P-Point is an ultrafilter such that for any partition of a set into subsets not in the ultrafilter, there exists a subset with finite intersections. A Q-Point is an ultrafilter where, for any partition of a set into finite pieces, there exists a subset with intersections of at most one element. P-Points and Q-Points are studied in fields such as topology [155, 305, 987, 880]. We explore P-Points and Q-Points within the context of an infinite connectivity system.
Definition 5.19.
Consider an infinite set $X$ equipped with a symmetric submodular function $f:2^{X}\to\mathbb{N}\cup\{\infty\}$. Suppose an infinite connectivity system $(X,f)$ satisfies $k$-Limit-Closed. Additionally, let $\{X_{n}\}_{n<\omega}$ be a symmetric submodular partition of $X$ such that $f(X_{n})\leq k$ for all $n$.
Ramsey (Selective) Point of Order $k+1$: A subset $A\subseteq X$ is called a Ramsey Point of order $k+1$ if for every symmetric submodular partition $\{X_{n}\}_{n<\omega}$ of $X$ such that $f(X_{n})\leq k$ and $X_{n}\notin F$ for all $n$, there exists an $A\in F$ such that:
$|A\cap X_{n}|\leq 1,\quad f(A\cap X_{n})\leq k,\quad\text{and}\quad f(A)\leq k\quad\text{for all }n<\omega.$
P-Point (Weakly Selective Point) of Order $k+1$: A subset $A\subseteq X$ is called a P-Point of order $k+1$ if for every symmetric submodular partition $\{X_{n}\}_{n<\omega}$ of $X$ such that $f(X_{n})\leq k$ and $X_{n}\notin F$ for all $n$, there exists an $A\in F$ such that:
$|A\cap X_{n}|<\omega,\quad f(A\cap X_{n})\leq k,\quad\text{and}\quad f(A)\leq k\quad\text{for all }n<\omega.$
Q-Point of Order $k+1$: A subset $A\subseteq X$ is called a Q-Point of order $k+1$ if for every symmetric submodular partition $\{X_{n}\}_{n<\omega}$ of $X$ into finite sets, such that $f(X_{n})\leq k$ and $X_{n}\notin F$ for all $n$, there exists an $A\in F$ such that:
$|A\cap X_{n}|\leq 1,\quad f(A\cap X_{n})\leq k,\quad\text{and}\quad f(A)\leq k\quad\text{for all }n<\omega.$
Semi-Q-Point (Rapid Point) of Order $k+1$: A subset $A\subseteq X$ is called a Semi-Q-Point of order $k+1$ if for every symmetric submodular partition $\{X_{n}\}_{n<\omega}$ of $X$ into finite sets, such that $f(X_{n})\leq k$ and $X_{n}\notin F$ for all $n$, there exists an $A\in F$ such that:
$|A\cap X_{n}|\leq n,\quad f(A\cap X_{n})\leq k,\quad\text{and}\quad f(A)\leq k\quad\text{for all }n<\omega.$
Theorem 5.20.
Consider an infinite set $X$ equipped with a symmetric submodular function $f:2^{X}\to\mathbb{N}\cup\{\infty\}$. Suppose an infinite connectivity system $(X,f)$ satisfies $k$-Limit-Closed. An ultrafilter $F$ on an infinite connectivity system $(X,f)$ is Ramsey (Selective) of order $k+1$ if and only if it is both a P-Point and a Q-Point of order $k+1$.
Proof 5.21.
Assume $F$ is a Ramsey ultrafilter of order $k+1$. By definition:
  1. 1.
    Since $F$ is Ramsey, for every symmetric submodular partition $\{X_{n}\}_{n<\omega}$ of $X$, there exists an $A\in F$ such that $|A\cap X_{n}|\leq 1$ and $f(A\cap X_{n})\leq k$ for all $n<\omega$.
    • This directly implies that $A$ also satisfies the conditions for a Q-Point, as it meets the finite intersection condition $|A\cap X_{n}|\leq 1$.
  2. 2.
    Similarly, the condition $|A\cap X_{n}|\leq 1$ (from being Ramsey) trivially implies $|A\cap X_{n}|<\omega$, satisfying the requirement for a P-Point.
Thus, every Ramsey ultrafilter of order $k+1$ is also a P-Point and a Q-Point of the same order.
Now, assume $F$ is both a P-Point and a Q-Point of order $k+1$. We need to show that $F$ is Ramsey of order $k+1$.
Let $\{X_{n}\}_{n<\omega}$ be a symmetric submodular partition of $X$ such that $X_{n}\notin F$ and $f(X_{n})\leq k$ for all $n$.
  1. 1.
    Since $F$ is a P-Point of order $k+1$, we can choose a set $A_{0}\in F$ such that $|A_{0}\cap X_{n}|<\omega$ and $f(A_{0}\cap X_{n})\leq k$ for all $n<\omega$.
  2. 2.
    Enumerate the elements of $X\setminus\bigcup_{n<\omega}(A_{0}\cap X_{n})=\{a_{n}\}_{n<\omega}$ and define new sets $Y_{2n}:=A_{0}\cap X_{n}$ and $Y_{2n+1}:=\{a_{n}\}$.
  3. 3.
    Since $F$ is a Q-Point of order $k+1$, we can choose a set $A_{1}\in F$ such that $|A_{1}\cap Y_{n}|\leq 1$ and $f(A_{1}\cap Y_{n})\leq k$ for all $n<\omega$.
  4. 4.
    Define $A:=A_{0}\cap A_{1}$. Then $A\in F$, and for each $n<\omega$, we have:
    $|A\cap X_{n}|=|(A_{0}\cap X_{n})\cap A_{1}|=|Y_{2n}\cap A_{1}|\leq 1$
    and $f(A\cap X_{n})\leq k$.
Since $A$ satisfies the condition $|A\cap X_{n}|\leq 1$ for all $n$, and $f(A)\leq k$, $F$ is a Ramsey ultrafilter of order $k+1$.
@@ 5.2 Ultrafilter on Countable Connectivity System
Game 5.22 (Ultrafilter Game on a Countable Connectivity System).
@@ 5.3 Elimination Games with an Invisible Robber on a Countable Connectivity System
Game 5.27 (Elimination Game with an Invisible Robber).
@@ 6.1.6 Examination for Directed proper-path-width
We will explore the concept of proper-path-width for graphs. Proper-path-width [1131, 1130, 1138, 1133] measures how "tightly" a graph can be decomposed into overlapping vertex subsets, ensuring that no two subsets are identical while maintaining specific connectivity constraints. We are interested in extending this definition to directed graphs. Although still in the conceptual stage, the definition is outlined below [485].
@@ 6.1.7 Fuzzy Tree-width
Example 6.17.
(cf.[224, 503, 519]) Consider a fuzzy graph $G=(\sigma,\mu)$ with four vertices $V=\{v_{1},v_{2},v_{3},v_{4}\}$, as depicted in the diagram.
The membership degrees of the vertices are as follows:
$\sigma(v_{1})=0.1,\quad\sigma(v_{2})=0.3,\quad\sigma(v_{3})=0.2,\quad\sigma(v_{4})=0.4$
The fuzzy relation on the edges is defined by the values of $\mu$, where $\mu(x,y)\leq\sigma(x)\land\sigma(y)$ for all $x,y\in V$. The fuzzy membership degrees of the edges are as follows:
$\mu(v_{1},v_{2})=0.1,\quad\mu(v_{2},v_{3})=0.1,\quad\mu(v_{3},v_{4})=0.1$
$\mu(v_{4},v_{1})=0.1,\quad\mu(v_{2},v_{4})=0.3$
In this case, the fuzzy graph $G$ has the following properties:
  • Vertices $v_{1},v_{2},v_{3},v_{4}$ are connected by edges with varying membership degrees.
  • The fuzzy relations ensure that $\mu(x,y)$ for any edge $(x,y)$ does not exceed the minimum membership of the corresponding vertices.
@@
Although still in the conceptual stage, the definitions are outlined below (cf.[502]).
@@
(cf.[502]) Let $\xi=(V,\sigma,\mu,E)$ be a directed fuzzy graph. A fuzzy directed tree-decomposition of $\xi$ is a triple $(T,X,W)$, where:
@@
(cf.[502]) Let $\xi=(V,\sigma,\mu,E)$ be a directed fuzzy graph. A fuzzy directed path-decomposition of $\xi$ is a sequence $(X_{1},\dots,X_{r})$ of fuzzy subsets of $V$, called bags, that satisfies the following conditions:
@@
(cf.[502]) Let $\xi=(V,\sigma,\mu,E)$ be a directed fuzzy graph. A layout of $\xi$ is a bijection $f:V\rightarrow\{1,2,\dots,|V|\}$. The fuzzy directed bandwidth of the layout $f$ is defined as:
@@ 6.1.8 Graph parameters for Fuzzy HyperGraph
In fuzzy graph theory, the concept of a fuzzy hypergraph is well-known. It extends the notion of a hypergraph by incorporating the concepts of fuzzy sets and fuzzy graphs. Numerous studies have explored its applications in real-world scenarios[915, 128, 35, 852, 826, 1193, 859, 854].
@@
We are interested in exploring whether new insights can be gained by extending the graph width parameter known as hypertree-width to fuzzy hypergraphs. Given the vast range of applications for fuzzy hypergraphs, such an extension could potentially accelerate the development of real-world applications[35, 854, 1180, 128, 1192, 276]. For further information on Tree-width and Width Parameters related to other real-world uncertainties, please refer to Appendix E.
@@
[502] Let $H=(\mu,\rho)$ be a fuzzy hypergraph defined on a non-empty set $X$, where:
@@ 6.1.9 Directed Hypertree Width
We are interested in exploring whether new insights can be gained by extending the graph width parameter known as hypertree-width to directed hypergraphs (cf.[1047]). We will also explore the concept of fuzzy directed hypertree-width. Given the active research on the real-world applications of fuzzy directed hypergraphs (ex.[33, 853, 922, 924, 923]), studying fuzzy directed hypertree-width is both meaningful and timely.
@@ 6.1.10 Weighted Width
We also plan to explore width parameters such as Weighted Directed Tree-width, Weighted Directed cut-width, Weighted Directed Tree-breadth, Weighted Directed Hypertree-width and Weighted Directed Path-width in the context of weighted digraphs. Given that weighted digraphs are practical and versatile graph classes with a wide range of applications[1001, 138, 918, 20, 591], defining these width parameters is both meaningful and valuable for various fields. Also we plan to explore about width-paremeters on Weighted hypergraph (cf.[1089, 888, 1084]).
@@ 6.1.11 Directed Linear-width
We will consider about Directed Linear-width[488]. Although still in the conceptual stage, the definitions are outlined below. This is the linear version of the Branch-width of a directed graph[231] and is also an extension of the Linear-width of an undirected graph[1141, 1147].
@@ 6.1.12 Directed Connectivity system
In the future, we aim to explore the results when extending to variants of directed graphs. Similar to undirected graphs, directed graphs have various proposed classes such as directed multigraphs[310], complete directed graphs[63], quasi-transitive digraphs [101], semicomplete multipartite digraphs[1238], and tournaments[103]. We plan to investigate how these results might apply when extended to these classes.
And the graph class known as Bidirected Graphs [200, 1227] has also been a subject of study. Since it generalizes the concept of Directed Graphs, we anticipate that research in this area will continue to grow. In fact, Bidirected Tree-width has already been defined in the literature [1209]. Therefore, defining parameters such as Bidirected Path-width (the path version of Bidirected Tree-width), Bidirected Cut-width, Bidirected Branch-width, Bidirected Linear-width, and Bidirected Connectivity system and exploring their relationships, would be highly meaningful. Additionally, extending these concepts to bidirected hypergraphs is another avenue worth exploring (cf.[1115]).
6.1.13 DAG-PathWidth and Related Parameter
A Directed Acyclic Graph (DAG) is a graph with directed edges and no cycles, commonly used in blockchains[655, 746, 742, 647]. While DAG-width is the most well-known width parameter related to DAGs[130, 102, 131, 957], the concept of DAG-pathwidth has recently been proposed[746, 742]. In light of this, we aim to explore obstructions to DAG-pathwidth and define various parameters on DAGs. Although still in the conceptual stage, the anticipated definitions are provided below.
Definition 6.43.
The DAG-cut-width of a directed acyclic graph (DAG) $D=(V,E)$ is defined as:
[…]
where:
  • $\Phi(D)$ is the set of all linear layouts (topological orderings) of $D$.
  • $L(i,\varphi,D)$ represents the set of vertices mapped to positions $\leq i$ by the linear layout $\varphi$.
  • $R(i,\varphi,D)$ represents the set of vertices mapped to positions $>i$ by the linear layout $\varphi$.
The DAG-cut-width measures the maximum number of edges that must be "cut" when partitioning the vertices in a given topological order. It quantifies the complexity of the DAG in terms of edge connectivity across different partitions of the vertex set.
Definition 6.44.
The DAG-band-width of a directed acyclic graph $D=(V,E)$ is defined as:
[…]
where:
  • $\mathcal{F}(D)$ is the set of all bijective mappings $f:V\rightarrow\{1,2,\dots,|V|\}$, representing different vertex orderings.
  • $f(u)$ and $f(v)$ are the positions of vertices $u$ and $v$ in the layout defined by $f$.
We also consider tree-width, path-width, cut-width, and other parameters on a directed acyclic hypergraph (DAH). A directed acyclic hypergraph is a generalization of a directed acyclic graph, where each hyperedge can have multiple tails and heads [994, 1070]. Since the concept of directed acyclic hypergraphs is applied in areas such as graphical modeling, extending these width parameters to DAHs is meaningful.
@@ 6.2.1 Simple Game on a Connectivity System
A simple voting game is a mathematical model of voting systems in which a coalition of agents wins if it belongs to a specified set of "winning" coalitions, subject to conditions like monotonicity (cf. [479, 913, 312, 1137]). These simple games model voting power and offer a generalized interpretation of the concept of "majority." As discussed in previous sections regarding the relationship between ultrafilters and simple voting systems, this game is well-known to be closely connected to ultrafilters.
@@ 6.2.5 Ultrafilter Width under Small Set Connectivity Expansion Hypothesis
We investigate ultrafilters in the context of the Small Set Connectivity Expansion Hypothesis problem within Parameterized Complexity.
Parameterized Complexity is a framework in computational complexity theory (cf.[975, 81, 1091]) that offers a more nuanced analysis of algorithmic problems than traditional complexity measures, which focus solely on input size. In this approach, each problem instance includes an additional value called a parameter—typically a non-negative integer—that captures some aspect of the input’s structure or complexity. This allows us to understand how computational difficulty depends not just on the input size but also on this specific parameter. Due to its significance, Parameterized Complexity has been extensively studied  [376, 374, 375, 317, 367].
Formally, a parameterized problem $P$ is a subset of $\Sigma^{*}\times\mathbb{N}_{0}$, where $\Sigma$ is a finite alphabet, $\Sigma^{*}$ denotes all finite strings over $\Sigma$, and $\mathbb{N}_{0}$ is the set of non-negative integers. Each instance is a pair $(x,k)$, where $x$ is the main input and $k$ is the parameter.
A problem is called fixed-parameter tractable (FPT) (cf.[1055, 373, 372]) if there exists an algorithm that decides whether $(x,k)\in P$ in time $f(k)\cdot|x|^{O(1)}$, where:
  • $f(k)$ is a computable function depending only on $k$,
  • $|x|$ is the size of the input $x$,
  • $O(1)$ indicates that the exponent is a constant independent of $k$.
In other words, for small values of $k$, the problem can be solved efficiently even if $|x|$ is large, because the computational complexity is primarily influenced by the parameter $k$.
The class XP consists of parameterized problems solvable in time $|x|^{f(k)}$, where $f(k)$ depends on $k$. For each fixed $k$, the problem is solvable in polynomial time, but the degree of the polynomial depends on $k$. Each parameterized problem can be viewed as a collection of $k$-slices, denoted $P_{k}=\{x\mid(x,k)\in P\}$, where the parameter $k$ is fixed. For problems in FPT, each $P_{k}$ can be solved in polynomial time with a degree independent of $k$. In XP, however, the degree may increase with $k$.
A problem is para-NP-hard (cf.[668]) or para-NP-complete (cf.[563]) if there exists at least one fixed $k$ such that $P_{k}$ is NP-hard or NP-complete. This means that even when $k$ is fixed, the problem remains as hard as the most difficult problems in NP. To further classify parameterized problems, the W-hierarchy (cf.[274]) is introduced, consisting of complexity classes between FPT and XP:
[…]
These inclusions are believed to be strict. Problems that are W[1]-hard are considered unlikely to be fixed-parameter tractable; that is, they probably cannot be solved in time $f(k)\cdot|x|^{O(1)}$.
In computational complexity theory, a computational hardness assumption posits that a specific problem cannot be solved efficiently. One such pivotal assumption is the Small Set Expansion Hypothesis (SSEH), introduced by Raghavendra and Steurer [1012], which plays a critical role in understanding the hardness of approximation algorithms. The SSEH asserts that it is NP-hard to distinguish between small sets of vertices in a graph with low edge expansion (few edges leaving the set) and those with high edge expansion (many edges leaving the set). This hypothesis has become fundamental in analyzing the complexity of various graph algorithms. Since its inception, the SSEH has led to numerous inapproximability results, including those documented in [694, 1217, 93, 1232, 241].